# Fisher information flow in artificial neural networks

### v1

@@ II.5 LFI maximization
We demonstrate the successful application of this algorithm with the following example: we draw data from a normal distribution and employ the LFI maximization algorithm to compute the FI when the estimated parameter is the standard deviation. This is a worst case scenario in the sense that the LFI of the initial data set vanishes. We thus expect that extensive optimization is required, compared to the case where the LFI is initially close to its maximum. We observe that the distributions we encounter in our experiment are mostly of the latter type, where very little optimization is required (see Appendix E). These concepts are illustrated in Fig. 3 that depicts the estimated LFI during the maximization algorithm as a function of $d^{\prime}-d$. Each data point is $50$ dimensional with independent entries, drawn from a normal distribution with zero mean and a standard deviation of $\sigma=1\pm\Delta\sigma$. The step size is $\Delta\sigma=0.05$ and the sample size is 100 000100\,000. In this simple case the FI $I=\frac{2N}{\sigma^{2}}=100$ can be calculated analytically. The non-linearity is described in Appendix E and the variance of the noise term is $\sigma_{\mathrm{noise}}=0.1$. Examples of such curves for simulated and experimental data are displayed in Fig. A3 and Fig. A4, respectively, in Appendix E.
@@ III Fisher Information Flow
Equipped with this tool, we can now investigate how the FI flows through an ANN. For this purpose, we train models on two different sets of data (see Appendix D for details about the data sets): The first sample we use is drawn from a numerically generated multivariate normal distribution where the parameter of interest is a location parameter, proportional to the mean. The data for the second model originates from the experiment illustrated in Fig. 1, where the network is trained to estimate the horizontal position of the “space invader” from the recorded camera images. Details about the model, the data generation, the calculation of FI and shortcomings of the LFI maximization algorithm are specified in Appendix D.
In Fig. 4, we show the FI as a function of the layer depth at different stages of training: panel (a) stands for the Gaussian data and panel (c) for the experimental data. The corresponding mean squared error (MSE) loss functions, normalized by the FI of the data set in the input layer (see Sec. IV) are shown in panel (b) for the Gaussian data and in panel (d) for the experimental data. The loss functions are evaluated using the training data and additionally, using an independent validation data set that was not processed by the network during its training.
At epoch 0, i.e., when the weights and biases are randomly initialised, a significant portion of FI is lost whenever a non invertible transformation acts upon the data. This behavior reflects the fact that inadequately adjusted weights and biases in the ANN manifest as sinks for the flow of FI through the network. However, after sufficiently many training epochs, a large portion of the FI reaches the output layer. As the network is trained, it approaches optimal performance, where the variance of the data in the output is minimal and as a consequence, the network conserves the FI of the data that is processed by the network.
@@
Figure 4: Fisher information calculated for a generic parameter (see Sec. D): (a,b) for numerically-generated data sampled from a Gaussian distribution and (c,d) for experimental data from the setup in Fig. 1. Panels (a) and (c) show the FI for the corresponding data set as a function of layer depth, normalized by the FI of the input data set. With increased training of the network (see color bars for the number of training epochs), the FI in all layers approaches the FI contained in the input data. Correspondingly, the normalized learning curves drop for an increasing number of training epochs: panels (b) and (d) show the training and validation loss (the MSE cost function estimated with the training data and the validation data), as a function of the epoch number.
@@
So far, we tracked the flow of FI evaluated at one specific parameter. However, a special feature of the FI is its locality, which means that the FI generally depends on the true parameter value. Interestingly, our method can be used to resolve the parameter dependence of the information flow itself. As we show in Appendix F such a parameter-resolved analysis indeed reveals non-trivial behavior that would remain hidden with a global analysis.
With our framework based on statistical estimation theory we are able to provide a quantitative characterization of propagation of information through the neural network throughout the training process, which is a significant step towards opening the black box of ANNs when they perform estimation tasks. Given the difficulty of estimation problems with arbitrary distributions of the data, the computational cost of our method is very low and we could perform the computations on a commercial CPU. Moreover, for both the Gaussian and the experimental training data set, the LFI itself is already close to the true FI in most of the ANNs layers, including the input layer. For these distributions the dimension in which the data is embedded does not need to be very large and the LFI maximisation algorithm converges quickly.
@@ IV Early stopping based on Fisher information
The ultimate precision limit in the estimation process is reached, when the ANN uses all the FI $I$ contained in the input data for the estimate provided in the output layer. From this reasoning we deduce that the mean square error (MSE) of the estimate that the ANN can optimally provide, is limited by the Cramér-Rao lower bound (CRLB):
@@
Most importantly, this relation tells us that our ability to quantify the FI for arbitrary distributions of the input data puts us in the unique position to formulate a criterion for early stopping using only the training data set, i.e., to stop training the network as soon as the condition in Eq. 7 is met. In this way our approach generalizes earlier studies [59, 60, 61, 62, 63], where the performance of ANNs was studied for very specific network architectures or under the assumption that the distribution of the input data is known or takes on a simple shape. Several other useful alternative strategies to conventional early stopping methods have also been developed by the deep learning community [64, 65, 66, 67, 68, 69]. What sets our method apart is that it is specifically applicable to ANNs trained to perform parameter estimation problems, and based on physical considerations.
@@
Specifically, a training loss larger or smaller than the inverse FI indicates under- or overfitting, respectively. This is illustrated in Fig. 5, where the MSE loss function is shown as function of the number of training epochs, both for the numerically generated training data (left panel) and the experimental data (right panel). In both cases, MSE loss functions are normalized by multiplying with the FI of the input data, so that for an estimator that achieves the CRLB the loss is equal to one.
@@
After the training loss in Fig. 5 crosses the horizontal line $\mathrm{MSE}\times I=1$, the validation loss is already very close to its global minimum and further training or hyperparameter tuning cannot improve the performance of the network. We observe that with this early stopping criterion, the epoch number of minimum validation loss is estimated remarkably well directly from the training loss. For the experimental data we observe that although the minimum of the normalized MSE at $\mathrm{MSE}\times I=1$ is not achieved precisely, the condition of stopping when the training loss falls below the CRLB still allows us to find a close approximation of the epoch where the validation loss is minimal.
@@
Note that if the complexity of the problem requires a much larger data set than what is available in order to perform near the CRLB, this criterion is not sufficient and the model requires validation. Indeed, we need to assume that the ANN does not perform too far off the CRLB for the dataset under study. If this condition is not satisfied, $\mathrm{MSE}\times I$ can reach unity when the network is already overfitting. Details about the neural networks and the numerical methods employed in this section are provided in Apps. D and E. Note that in Eq. (7) we assume that the FI of the input data does not vary with $\theta$, otherwise the inverse FI must be averaged over $\theta$. We also note that if a cost function other than the MSE is used, one needs to evaluate the CRLB, i.e., to compare the right hand side of Eq. (2) with the variance of the network’s output, in addition to the cost function. Moreover, for Eq. (7) to hold, the network must represent an unbiased estimator. If one expects a significant bias to occur, instead of Eq. (6) the generalized expression [22] $\sigma^{2}\geq(\partial_{\theta}\langle\hat{\theta}\rangle)^{2}/I$ should be used as the bound for the estimator’s variance. The numerator $(\partial_{\theta}\langle\hat{\theta}\rangle)^{2}$ in the equation above is different from $1$ if the expectation value of the estimated parameter $\langle\hat{\theta}\rangle$ does not match the true parameter, i.e., when the estimator is biased. For the networks and data sets we consider here, the assumptions of negligible bias and constant FI hold to a good approximation, thus Eq. (6) applies.
@@
Figure 5: Normalized MSE as a function of number of epochs for (a) numerically-generated training data with Gaussian noise and (b) data measured in the experiment. The dark blue/red lines denote the validation losses and the light blue/red lines depict the training losses. The loss function is chosen as the MSE normalized by multiplication with the FI of the input data ($\mathrm{MSE}\times I$). The dashed horizontal lines indicate where the loss is equal to $1$. The epoch number where the training loss crosses this horizontal line is marked by a vertical dashed line. This line coincides with the epoch number where the validation loss is near its minimum.
@@ V Discussion
With the methods introduced here, we are able to track the FI flow through an ANN. Based on this FI analysis, we observe how an ANN learns to preserve the information contained in high-dimensional image data while compressing it to a single dimension. Our approach does not require strong assumptions about the neural network architecture or the exact shape of the data distribution. As previously mentioned, a feature of our FI analysis is that it is local with respect to the parameter value. However, bounding the performance of the network, e.g., for early stopping based on training data only, requires a global measure of information. In this case, one either assumes that the FI of the training data is approximately independent of the parameter, which is the case for our data sets; alternatively if this assumption does not hold, one must evaluate the FI of the training data at multiple parameters, which means that the computational cost for calculating the FI for a single parameter is multiplied with the number of parameters where the FI is evaluated. For the performance bound, the FI is averaged over the range of possible parameter values, hence one can, instead of using a fine parameter-grid, also sample random parameters, evaluate the FI at each random parameter and sum them up in order to approximate this average.
@@
In order to illustrate that our approach is applicable to very different noise statistics, we employ in Appendix G ANNs that are trained to predict a parameter from a correlated Lognormal distribution. We track the flow of FI through an ANN and demonstrate that early stopping based on the training loss works very well also for this data set.
@@
There is a large number of degrees of freedom for choosing the shape of the non-linear transformations and the optimisation procedure for the LFI maximization algorithm. Further research in this direction might yield more efficient techniques for estimating the FI from data. In contrast to methods for calculating the FI that entail estimating the probability density function, our approach is data driven in the sense that it profits from the size of the data set, rather than from the simplicity of the distribution. A prerequisite for our approach to work is that the data set consists of a sufficiently large number of data points, so that the mean and the covariance matrix can be estimated accurately. Knowing the variance of our estimator of the FI (see Appendix A.2) and how it is affected by the choice of algorithmic parameters such as the sample size, allows us to verify whether these prerequisites are fulfilled. This insight also allows us to work out the boundaries of our algorithm and potentially also explore the regime of small data sets.
@@
Recently, non-parametric methods [70, 71, 72] for estimating the FI from a finite sized data set that do not rely on an estimate of the probability density, have been developed. For these approaches, deriving confidence bounds is very difficult. While the accuracy of our approach cannot be predicted if the FI is unknown [53], it has the advantage that for sufficiently large data sets it always gives a pessimistic approximation of the FI [51]. Such an estimate of a conservative replacement of the FI is valuable, as it does not only act as a close approximation but also yields a value for the precision that can be, at least locally (see Subsec. II.4), achieved by an estimate when the data is further processed [51]. Moreover, these methods typically involve functions that depend on all possible pairs of data points, i.e., $N_{\mathrm{sample}}^{2}$ where $N_{\mathrm{sample}}$ is the sample size, making them prohibitively expensive when a large data set is required. The LFI maximization method involves constructing covariance matrices with a number of entries given by the squared dimension of the space in which we project the data. However, the squared dimension is in our regime of large data sets much smaller than $N_{\mathrm{sample}}^{2}$, which makes our LFI maximization method less expensive in terms of memory. Another drawback of these non-parametric methods is their tendency to be very sensitive with respect to the choice of the spacing between the discrete values of the parameter in the data set, when evaluating the FI with a finite difference scheme [73]. We observe that our method yields inaccurate results when the spacing is very small, however, it yields close pessimistic approximations of the FI for larger spacings.
@@ VI Summary and outlook
Our work demonstrates how ANNs that are trained for performing estimation tasks, optimize their performance by maximizing the flow of FI from the input to the output layer. When all of the FI stored in the input data reaches the output layer, where the estimate is produced, the optimal performance at the CRLB is achieved. Tests on numerically generated data yield perfect agreement with this prediction. For the case of a concrete optical experiment, involving the estimation of an object’s position from noisy image data, we find that the trained network is able to channel around 80% of FI to the output layer and thus comes close to the CRLB although not exactly reaching it, which we discuss in Appendix A.1.
The non-trivial task of calculating the FI in the hidden layers is accomplished by using the quantity of LFI, which often is a close approximation of the FI and comes with the advantage of being easy to compute. In order to deal with situations where this approximation is insufficient, we present an algorithm that gradually improves upon the LFI by applying non-linear embeddings into spaces of ever-increasing dimensionalities, until convergence is achieved. This computational tool not only allows us to study the FI flow but also yields a criterion for early stopping, which only requires knowledge about the FI of the input data. From a conceptual point of view, our results demonstrate that similar to the information-conserving flow of FI mediated by electromagnetic fields [29], FI can also flow through suitably trained ANNs without any information getting lost along the way.
The opportunities for a FI flow analysis in ANNs are numerous, such as for identifying optimized network architectures, and for reducing the need for expensive trial and error approaches. Moreover, conservation of FI yields a condition for constructing an ANN layer-by-layer by choosing the weights such that the FI of one layer equals that of a previous layer. This could help to resolve problems such as vanishing/exploding gradients since one does not need to propagate the gradients through the whole network. Conservation of FI can also be used as a condition to compress the data to a smaller dimension, without losing information about the parameter of interest [58]. As a next step, it would also be interesting to include prior information of the data about the estimated parameter. Extending our analysis to networks with skip connections such as the popular ResNet architecture [74], where the layers of the network do not show the Markov chain structure (on which we relied when assuming that the FI flow obeys the data processing inequality), would be a natural next step towards opening the black box of more complex ANN models. Moreover, we aim to apply our FI analysis to physical neural networks [75], in order to gain a better understanding of how they learn to process information during their training [76, 77].

### v2

@@ II.5 LFI maximization
We demonstrate the successful application of this algorithm with the following example: We draw data from a normal distribution and employ the LFI maximization algorithm to compute the FI when the estimated parameter is the standard deviation. This is a worst-case scenario in the sense that the LFI of the initial dataset vanishes. We, thus, expect that extensive optimization is required, compared to the case where the LFI is initially close to its maximum. We observe that the distributions we encounter in our experiment are mostly of the latter type, where very little optimization is required (see Appendix E). These concepts are illustrated in Fig. 3 that depicts the estimated LFI during the maximization algorithm as a function of $d^{\prime}-d$. Each data point is $50$ dimensional with independent entries, drawn from a normal distribution with zero mean and a standard deviation of $\sigma=1\pm\Delta\sigma$. The step size is $\Delta\sigma=0.05$ and the sample size is 100 000100\,000. In this simple case, the FI for the estimation of the mean $I=\frac{2N}{\sigma^{2}}=100$ can be calculated analytically. The nonlinearity is described in Appendix E and the variance of the noise term is $\sigma_{\mathrm{noise}}=0.1$. Examples of such curves for simulated and experimental data are displayed in Figs. A3 and A4, respectively, in Appendix E.
@@ III Fisher Information Flow
Equipped with this tool, we can now investigate how the FI flows through an ANN. For this purpose, we train models on two different sets of data (see Appendix D for details about the datasets): The first sample we use is drawn from a numerically generated multivariate normal distribution where the parameter of interest is a location parameter, proportional to the mean. The data for the second model originate from the experiment illustrated in Fig. 1, where the network is trained to estimate the horizontal position of the “space invader” from the recorded camera images. Details about the model, the data generation, the calculation of FI, and shortcomings of the LFI maximization algorithm are specified in Appendix D.
In Fig. 4, we show the FI as a function of the layer depth at different stages of training: Figure 4(a) stands for the Gaussian data and Fig. 4(c) for the experimental data. The corresponding mean squared error (MSE) loss functions, normalized by the FI of the dataset in the input layer (see Sec. IV), are shown in Fig. 4(b) for the Gaussian data and in Fig. 4(d) for the experimental data. The loss functions are evaluated using the training data and additionally, using an independent validation dataset that was not processed by the network during its training.
At epoch 0, i.e., when the weights and biases are randomly initialized, a significant portion of FI is lost whenever a noninvertible transformation acts upon the data. This behavior reflects the fact that inadequately adjusted weights and biases in the ANN manifest as sinks for the flow of FI through the network. However, after sufficiently many training epochs, a large portion of the FI reaches the output layer. As the network is trained, it approaches optimal performance, where the variance of the data in the output is minimal and, as a consequence, the network conserves the FI of the data that is processed by the network.
@@
Figure 4: Fisher information calculated for a generic parameter (see Appendix D): (a),(b) for numerically generated data sampled from a Gaussian distribution and (c),(d) for experimental data from the setup in Fig. 1. Panels (a) and (c) show the FI for the corresponding dataset as a function of layer depth, normalized by the FI of the input dataset. With increased training of the network (see color bars for the number of training epochs), the FI in all layers approaches the FI contained in the input data. Correspondingly, the normalized learning curves drop for an increasing number of training epochs: panels (b) and (d) show the training and validation loss (the MSE cost function estimated with the training data and the validation data) as a function of the epoch number.
@@
So far, we tracked the flow of FI evaluated at one specific parameter. However, a special feature of the FI is its locality, which means that the FI generally depends on the true parameter value. Interestingly, our method can be used to resolve the parameter dependence of the information flow itself. As we show in Appendix F, such a parameter-resolved analysis indeed reveals nontrivial behavior that would remain hidden with a global analysis.
With our framework based on statistical estimation theory we are able to provide a quantitative characterization of propagation of information through the neural network throughout the training process, which is a significant step toward opening the black box of ANNs when they perform estimation tasks. Given the difficulty of estimation problems with arbitrary distributions of the data, the computational cost of our method is very low and we could perform the computations on a commercial CPU. Moreover, for both the Gaussian and the experimental training datasets, the LFI itself is already close to the true FI in most of the ANNs layers, including the input layer. For these distributions the dimension in which the data are embedded does not need to be very large and the LFI maximization algorithm converges quickly.
@@ IV Early stopping based on Fisher information
The ultimate precision limit in the estimation process is reached, when the ANN uses all the FI $I$ contained in the input data for the estimate provided in the output layer. From this reasoning we deduce that the MSE of the estimate that the ANN can optimally provide, is limited by the CRLB:
@@
Most importantly, this relation tells us that our ability to quantify the FI for arbitrary distributions of the input data puts us in the unique position to formulate a criterion for early stopping using only the training dataset, i.e., to stop training the network as soon as the condition in Eq. (7) is met. In this way, our approach generalizes earlier studies [59, 60, 61, 62, 63], where the performance of ANNs was studied for very specific network architectures or under the assumption that the distribution of the input data is known or takes on a simple shape. Several other useful alternative strategies to conventional early stopping methods have also been developed by the deep learning community [64, 65, 66, 67, 68, 69]. What sets our method apart is that it is specifically applicable to ANNs trained to perform parameter estimation problems, and based on physical considerations.
@@
Specifically, a training loss larger or smaller than the inverse FI indicates under- or overfitting, respectively. This is illustrated in Fig. 5, where the MSE loss function is shown as function of the number of training epochs, for both the numerically generated training data (left) and the experimental data (right). In both cases, MSE loss functions are normalized by multiplying with the FI of the input data, so that for an estimator that achieves the CRLB the loss is equal to one.
@@
After the training loss in Fig. 5 crosses the horizontal line $\mathrm{MSE}\times I=1$, the validation loss is already very close to its global minimum and further training or hyperparameter tuning cannot improve the performance of the network. We observe that with this early stopping criterion, the epoch number of minimum validation loss is estimated remarkably well directly from the training loss. For the experimental data, we observe that, although the minimum of the normalized MSE at $\mathrm{MSE}\times I=1$ is not achieved precisely, the condition of stopping when the training loss falls below the CRLB still allows us to find a close approximation of the epoch where the validation loss is minimal.
@@
Note that if the complexity of the problem requires a much larger dataset than what is available in order to perform near the CRLB, this criterion is not sufficient and the model requires validation. Indeed, we need to assume that the ANN does not perform too far off the CRLB for the dataset under study. If this condition is not satisfied, $\mathrm{MSE}\times I$ can reach unity when the network is already overfitting. Details about the neural networks and the numerical methods employed in this section are provided in Appendixes D and E. Note that in Eq. (7) we assume that the FI of the input data does not vary with $\theta$; otherwise, the inverse FI must be averaged over $\theta$. We also note that if a cost function other than the MSE is used, one needs to evaluate the CRLB, i.e., to compare the right-hand side of Eq. (2) with the variance of the network’s output, in addition to the cost function. Moreover, for Eq. (7) to hold, the network must represent an unbiased estimator. If one expects a significant bias to occur, instead of Eq. (6), the generalized expression [22] $\sigma^{2}\geq(\partial_{\theta}\langle\hat{\theta}\rangle)^{2}/I$ should be used as the bound for the estimator’s variance. The numerator $(\partial_{\theta}\langle\hat{\theta}\rangle)^{2}$ in the equation above is different from $1$ if the expectation value of the estimated parameter $\langle\hat{\theta}\rangle$ does not match the true parameter, i.e., when the estimator is biased. For the networks and datasets we consider here, the assumptions of negligible bias and constant FI hold to a good approximation; thus, Eq. (6) applies.
@@
Figure 5: Normalized MSE as a function of number of epochs for (a) numerically generated training data with Gaussian noise and (b) data measured in the experiment. The dark blue and red lines denote the validation losses, and the light blue and red lines depict the training losses. The loss function is chosen as the MSE normalized by multiplication with the FI of the input data ($\mathrm{MSE}\times I$). The dashed horizontal lines indicate where the loss is equal to $1$. The epoch number where the training loss crosses this horizontal line is marked by a vertical dashed line. This line coincides with the epoch number where the validation loss is near its minimum.
@@ V Discussion
With the methods introduced here, we are able to track the FI flow through an ANN. Based on this FI analysis, we observe how an ANN learns to preserve the information contained in high-dimensional image data while compressing it to a single dimension. Our approach does not require strong assumptions about the neural network architecture or the exact shape of the data distribution. A feature of our FI analysis is that it is local with respect to the parameter value. However, bounding the performance of the network, e.g., for early stopping based on training data only, requires a global measure of information. In this case, one either assumes that the FI of the training data is approximately independent of the parameter, which is the case for our datasets; alternatively, if this assumption does not hold, one needs to evaluate the FI of the training data at multiple parameters, which means that the computational cost for calculating the FI for a single parameter is multiplied with the number of parameters where the FI is evaluated. For the performance bound, the FI is averaged over the range of possible parameter values; hence, one can, instead of using a fine parameter grid, also sample random parameters, evaluate the FI at each random parameter and sum them up in order to approximate this average.
@@
In order to illustrate that our approach is applicable to very different noise statistics, we employ in Appendix G ANNs that are trained to predict a parameter from a correlated log-normal distribution. We track the flow of FI through an ANN and demonstrate that early stopping based on the training loss works very well also for this dataset.
@@
There is a large number of degrees of freedom for choosing the shape of the nonlinear transformations and the optimization procedure for the LFI maximization algorithm. Further research in this direction might yield more efficient techniques for estimating the FI from data. In contrast to methods for calculating the FI that entail estimating the probability density function, our approach is data driven in the sense that it profits from the size of the dataset rather than from the simplicity of the distribution. A prerequisite for our approach to work is that the dataset consists of a sufficiently large number of data points, so that the mean and the covariance matrix can be estimated accurately. Knowing the variance of our estimator of the FI (see Appendix A.2) and how it is affected by the choice of algorithmic parameters such as the sample size, allows us to verify whether these prerequisites are fulfilled. This insight also allows us to work out the boundaries of our algorithm and potentially also explore the regime of small datasets.
@@
Recently, nonparametric methods [70, 71, 72] for estimating the FI from a finite-sized dataset that do not rely on an estimate of the probability density, have been developed. For these approaches, deriving confidence bounds is very difficult. While the accuracy of our approach cannot be predicted if the FI is unknown [53], it has the advantage that for sufficiently large datasets it always gives a pessimistic approximation of the FI [51]. Such an estimate of a conservative replacement of the FI is valuable, as it not only acts as a close approximation but also yields a value for the precision that can be, at least locally (see Sec. II.4), achieved by an estimate when the data are further processed [51]. Moreover, these methods typically involve functions that depend on all possible pairs of data points, i.e., $N_{\mathrm{sample}}^{2}$, where $N_{\mathrm{sample}}$ is the sample size, making them prohibitively expensive when a large dataset is required. The LFI maximization method involves constructing covariance matrices with a number of entries given by the squared dimension of the space in which we project the data. However, the squared dimension is in our regime of large datasets much smaller than $N_{\mathrm{sample}}^{2}$, which makes our LFI maximization method less expensive in terms of memory. Another drawback of these nonparametric methods is their tendency to be very sensitive with respect to the choice of the spacing between the discrete values of the parameter in the dataset, when evaluating the FI with a finite-difference scheme [73]. We observe that our method yields inaccurate results when the spacing is very small; however, it yields close pessimistic approximations of the FI for larger spacings.
@@ VI Summary and outlook
Our work demonstrates how ANNs that are trained for performing estimation tasks, optimize their performance by maximizing the flow of FI from the input to the output layer. Using numerically generated data, we observe that when all of the FI stored in the input data reaches the output layer, where the estimate is produced, the optimal performance at the CRLB is achieved. For the case of a concrete optical experiment, involving the estimation of an object’s position from noisy image data, we find that the trained network is able to channel around 80% of FI to the output layer and, thus, comes close to the CRLB although not exactly reaching it, which we discuss in Appendix A.1.
The nontrivial task of calculating the FI in the hidden layers is accomplished by using the quantity of LFI, which often is a close approximation of the FI and comes with the advantage of being easy to compute. In order to deal with situations where this approximation is insufficient, we present an algorithm that gradually improves upon the LFI by applying nonlinear embeddings into spaces of ever-increasing dimensionalities, until convergence is achieved. This computational tool not only allows us to study the FI flow, but also yields a criterion for early stopping, which requires only knowledge about the FI of the input data. From a conceptual point of view, our results demonstrate that, similar to the information-conserving flow of FI mediated by electromagnetic fields [29], FI can also flow through suitably trained ANNs without any information getting lost along the way.
The opportunities for a FI flow analysis in ANNs are numerous, such as for identifying optimized network architectures, and for reducing the need for expensive trial and error approaches. Moreover, conservation of FI yields a condition for constructing an ANN layer by layer by choosing the weights such that the FI of one layer equals that of a previous layer. This could help to resolve problems such as vanishing and exploding gradients, since one does not need to propagate the gradients through the whole network. Conservation of FI can also be used as a condition to compress the data to a smaller dimension, without losing information about the parameter of interest [58]. As a next step, it would also be interesting to include prior information of the data about the estimated parameter. Extending our analysis to networks with skip connections such as the popular ResNet architecture [74], where the layers of the network do not show the Markov chain structure (on which we relied when assuming that the FI flow obeys the data processing inequality), would be a natural next step toward opening the black box of more complex ANN models. Moreover, we aim to apply our FI analysis to physical neural networks [75], in order to gain a better understanding of how they learn to process information during their training [76, 77].